How many 5-cards hands having exactly 3 aces and 2 other cards can be dealt?
Answer by stanbon(75887)

How many 5-cards hands having exactly 3 aces and 2 other cards can be dealt?
-----------
Pick 3 aces: 4C3 = 4 ways
Pick 2 other cards: 48C2 = 48*47/1*2 = 1128 ways
----------------------------------
Ans: 4*1128 = 4512 ways
=========================
Cheers,
Stan H.